OpenStudy (anonymous):

A ball is hurled in the air such that its horizontal (x) and vertical (y) shifts are given by the functions x(t) = 20t and y(t) = 40t − 5t2, where x and y are in meters and the time t is in seconds. What is the ratio of y to x at time t = 5 seconds?

OpenStudy (lochana):

at time t =5 x(5) = 20*5 = 100m ; y(5)= 40*5- 5(5)^2( assumed ^ was missed in Q) , so y(5)= 200-5*25 = 200-125 = 75m thus ratio between x and y is x/y= 100/75= 4/3